### v1

@@ Abstract
In cases with multi-antenna users, the APs can either transmit the same or different data streams. The latter reduces the fronthaul signaling but comes with a SE loss. We propose precoding and combining schemes for these cases and consider whether channel knowledge is shared between the APs. Finally, we show numerically how the number of users, APs, and the number of antennas on users and APs affect the SE.
Index Terms:
Cell-free massive MIMO, multi-antenna users, downlink pilots, spectral efficiency, stream allocation.
@@ I Introduction
Cell-free massive multiple-input multiple-output (MIMO) is a wireless communication paradigm that has attracted great interest due to the vision of delivering uniformly high spectral efficiency (SE) over the coverage area [1, 2, 3]. In cell-free massive MIMO, a large number of geographically distributed access points (APs) cooperate to serve the users on the same time-frequency resources through coherent joint transmission where interference is managed by MIMO methods [4].
@@
The recent surveys [3, 2] highlight a large body of work on cell-free massive MIMO systems. The predominant assumption is single-antenna users, although user devices (e.g., phones and tablets) have been equipped with multiple antennas since 4G. Four antennas are the current norm for mid-band 5G devices, and larger antenna numbers are expected in 6G devices operating at higher frequencies. The few papers on cell-free massive MIMO that consider multi-antenna users show SE improvements (e.g., [5]) but assume perfect channel state information (CSI). A detailed SE analysis with realistic imperfect CSI is missing, particularly for downlink operation.
@@
The uplink SE with zero-forcing (ZF) combining and multi-antenna users was studied in [6] under the assumption of independent and identically distributed (i.i.d.) Rayleigh fading. This analysis was extended to consider the Weichselberger channel model in [7], where four operation regimes were compared. The uplink was further studied in [8, 9] with a focus on distortion caused by low-resolution hardware. The downlink performance with ZF precoding is analyzed in [10] under i.i.d. Rayleigh fading, using the hardening bound where the user device lacks CSI. Moreover, the downlink performance with maximum ratio (MR) precoder and downlink channel estimation was studied in [11]. This work was extended in [12] to consider the downlink with low-resolution digital-to-analog converters (DACs). Multi-antenna users have also been studied in the context of cellular massive MIMO. A single-cell scenario was considered in [13] and the hardening bound was used to compute an achievable downlink SE without CSI at the users. [14] computed a closed-form uplink SE expression in single-cell massive MIMO systems with multi-antenna users, MR combining, and Rician fading. Downlink MR precoding with multi-antenna users was considered in [15], with a focus on making the hardening bound tighter with an unusual precoding normalization. The correlation between the channel vectors to different antennas at the same user was studied in [16], and it was shown that asymptotic orthogonality cannot be guaranteed. Synchronization issues, such as phase misalignments between APs, were studied in [17] by proposing transmission methods and optimization of the AP clusters.
@@
In [18], it was shown that no downlink pilots are needed in massive MIMO systems with single-antenna users. The receiving users can perform decoding properly without CSI since the effective channel’s phase is removed by precoding, and the gain approaches its mean value—a phenomenon called channel hardening. The so-called hardening bound on the capacity gives a performance close to the perfect CSI case also in cell-free massive MIMO [4] with the possible exception of scenarios with MR precoding, if the precoding vectors are normalized incorrectly [19, 20] and when the effective channel means are normalized with their estimates [21]. In cases with very limited channel hardening (e.g., keyhole channels), higher SE can be achieved by explicitly estimating the effective channel’s amplitude from the received downlink signals [18]. These single-antenna results might be the reason why the hardening bound has also been used in [10, 12, 13, 15] when considering multi-antenna users. However, we will show that the use of downlink pilots can lead to great SE improvements in these cases.
@@ I-A Motivation and Contributions
In this paper, we analyze the performance of a cell-free massive MIMO system with multiple antennas on the users and APs, in a scenario with correlated Rayleigh fading channels. Conventional cell-free massive MIMO builds on the assumption that the serving APs send the same data stream to the user. However, we can send multiple streams to multi-antenna users, which enables an alternative transmission method where the serving APs send different streams to the same user non-coherently. The latter reduces the fronthaul signaling since only parts of the downlink data are sent to each AP. We analyze both transmission methods with arbitrary precoding to derive general SE expressions. In particular, we derive SE expressions for the cases when the user has CSI obtained through pilot-based downlink channel estimation.
We propose new precoders for the two transmission methods. These are needed because the conventional precoding schemes for multi-antenna users in [22, 23, 24] are not designed for situations with per-AP power constraints, partial CSI, and separate stream transmission. We further investigate different combining methods that suppress interference using the downlink channel estimates.
@@
In the conference version of this work [25], we assumed i.i.d. Rayleigh fading, centralized operation, and ZF combining and focused on showing the necessity of downlink pilots. In this paper, we further analyze multi-antenna users in cell-free massive MIMO by assuming correlated Rayleigh fading, extending the system model, and exploring new transmission and receive processing methods. We propose two different transmission methods for multi-antenna users. We propose precoders and derive the SE relations for the general combiner case. Moreover, we propose an MMSE combiner at the downlink and show the improvement over the ZF combiner.
The rest of this paper is organized as follows. In Section II, we describe the channel model and estimation of the channel in the uplink. In Section II-B, downlink data transmission methods are explained for both same and separate stream methods. In Section III, achievable SE results for the same stream transmission method are derived, particularly with downlink pilots. In Section IV, SE results are extended to the case with separate stream transmission. Next, in Section V, precoders are proposed for all the considered transmission methods. We explain under which channel conditions these methods coincide. Section VI provides numerical results where we compare the performance of the considered methods and compare different SE expressions. Finally, conclusions and future directions are given in Section VII.
Notation: The superscripts $(\cdot)^{*}$$(\cdot)^{\intercal}$ and $(\cdot)^{\mathrm{H}}$ denote the conjugate, transpose and conjugate transpose, respectively. Column vectors and matrices are denoted with boldface lowercase $\mathbf{x}$ and uppercase $\mathbf{X}$ letters. The expectation operation, variance, and covariance of a vector and matrix are denoted as $\mathbb{E}\{\cdot\},\mathrm{var}(\cdot),\mathrm{cov}(\cdot)$. The covariances are evaluated as $\mathbf{C}_{\mathbf{a}\mathbf{b}}=\mathbb{E}\{\mathbf{a}\mathbf{b}^{\mathrm{H}}\}$ for $\mathbf{a}$ and $\mathbf{b}$ vectors. The trace and determinant of a matrix are given by $\mathrm{tr}(\cdot),|\cdot|$. Convergence is shown with $\rightarrow$. The $L_{2}$ norm of a vector is given by $\|\mathbf{x}\|_{2}$ and the Frobenius norm of a matrix is given by $\|\mathbf{X}\|_{F}$. The $N$-dimensional identity matrix is denoted by $\mathbf{I}_{N}$. $\mathbf{x}\sim\mathcal{N}_{\mathbb{C}}(\mathbf{0},\mathbf{R})$ denotes a circularly symmetric complex Gaussian random vector $\mathbf{x}$ with covariance matrix $\mathbf{R}$. The vectorization of a matrix $\mathbf{X}$ is denoted by $\mathrm{vec}(\mathbf{X})$.
@@ II System Model
We consider a cell-free massive MIMO system with $L$ APs and $K$ users arbitrarily distributed in a large geographical area. Each AP has $N$ antennas and each user has $M$ antennas. We consider the standard block-fading time-division duplexing (TDD) operation where the channel between AP $l$ and user $k$ in an arbitrary coherence block is denoted by $\mathbf{H}_{lk}\in\mathbb{C}^{N\times M}$ [4]. We assume a correlated Rayleigh fading channel model. Hence, the channel between AP $l$ and user $k$ is complex Gaussian distributed with the spatial correlation matrix $\mathbf{R}_{lk}\in\mathbb{C}^{NM\times NM}$, which is defined for the vectorization $\mathrm{vec}(\mathbf{H}_{lk})$ of the channel matrix $\mathbf{H}_{lk}$. The channel covariance is expressed through the covariances of its individual column vectors. To obtain this covariance, we vectorize the channel matrix.
@@ II-A Uplink Channel Estimation
As customary in TDD operation, the APs estimate the channels in the uplink and use the estimates in both uplink data reception and downlink data transmission [4]. In this paper, we will focus on uplink channel estimation and downlink data transmission. All users send uplink pilot sequences, and the APs estimate their channels to the users based on their received signals. Each user sends a $\tau_{\mathrm{p}}$-length orthonormal pilot sequence per transmit antenna. We will let $\tau_{\mathrm{c}}$ denote the length of the coherence block (in the number of symbols); thus, the pilot signaling overhead is $\tau_{\mathrm{p}}/\tau_{\mathrm{c}}$.
@@
$=\sqrt{q_{k}\tau_{\mathrm{p}}}\mathrm{vec}(\mathbf{H}_{lk})+\sum_{{i\in\mathcal{P}_{k}\setminus\{k\}}}\sqrt{q_{i}\tau_{\mathrm{p}}}\mathrm{vec}(\mathbf{H}_{li})+\mathrm{vec}(\mathbf{N}_{lk})$
@@
where we introduced the notation $\mathbf{h}_{li}=\mathrm{vec}(\mathbf{H}_{li})$ for the vectorization of the channel matrix and $\mathbf{n}_{lk}=\mathrm{vec}(\mathbf{N}_{lk})$ for the noise. The MMSE estimate of $\mathbf{h}_{lk}$ is [26, Ch. 3]
@@ II-B1 Downlink with Same Stream Transmission
The most advanced downlink operation involves all APs in the transmission of all data streams, using the centralized operation [4]. Hence, the APs cooperate in the transmission, share CSI obtained from the uplink channel estimation, and precode their transmitted downlink signals based on it.
@@
For a given realization of the effective channels, created by the channel realizations and precoding scheme, the achievable SE varies depending on the available CSI at the receiver. We will provide expressions for three different cases in Section III and compare them numerically in Section VI.
@@ III-A Hardening Bound without CSI at the Receiver
When the receiver lacks CSI regarding the instantaneous channel realizations, the SE can be lower bounded using the hardening bound [4]. This method is conventionally used for single-antenna users, but the multi-antenna extension was considered in [10, 12, 13, 11, 7, 15]. The main idea is that user $k$ knows the statistics of its effective channel $\mathbf{B}_{kk}=\mathbf{H}_{k}^{\rm{H}}\mathbf{W}_{k}$, such as its mean $\bar{\mathbf{B}}_{kk}=\mathbb{E}\left\{\mathbf{B}_{kk}\right\}$, and the precoder matrix has been selected based on the CSI to make the variations around the mean small. We rewrite the received signal in (II-B1) as
@@
where ||$\cdot$|| ⋅ | denotes the determinant.
The proof of this lemma builds on interpreting (8) as a point-to-point MIMO system with the deterministic channel matrix $\bar{\mathbf{B}}_{kk}$ and uncorrelated additive noise with the covariance matrix $\mathbf{\Xi}_{k}\in\mathbb{C}^{M\times M}$, which can be obtain by using sample averages. The SE expression is then obtained as a lower bound on the channel capacity by utilizing the worst-case uncorrelated additive noise term [27]. The single-antenna version of Lemma 1 is called the hardening bound [4] since the effective channel realization $\mathbf{B}_{kk}$ is close to its mean $\bar{\mathbf{B}}_{kk}$ (in relative terms) when $M=1$, $N$ is large, and the precoder is selected as MR, ZF, or a filter of similar structure. This phenomenon is called channel hardening. However, as the effective channel is a matrix instead of a scalar when $M>1$, the same kind of channel hardening is not achieved herein since each column of the precoder matrix can only achieve hardening with respect to one row in the channel matrix. The resulting performance loss is demonstrated later in this paper.
@@
We conclude that only the diagonal entries of $\mathbf{B}_{kk}$ harden, while the off-diagonal entries that represent the inter-stream interference do not. There are no off-diagonal entries when $M=1$, so channel hardening holds in that special case (as shown in [18, 4]), while most entries do not harden when $M>1$. Therefore, the hardening bound is loose when $M>1$, which calls for developing a methodology for acquiring CSI at the receiving user using downlink pilots. [^1] The pilot-free downlink CSI acquisition method proposed in [18] cannot be extended to the multi-antenna case because it relies on the proposal that the precoder matrix is selected to make the entries of $\mathbf{B}$ (approximately) real-valued and positive. This can only be guaranteed for $M$ entries of $\mathbf{B}$, while the remaining $M^{2}-M$ (typically off-diagonal entries) cannot be estimated in this way.

### v2

@@ Abstract
In cases with multi-antenna users, the APs can either coherently transmit the same data streams, or alternatively they transmit separate data streams non-coherently. The latter approach reduces the fronthaul signaling overhead, albeit with a potential SE penalty. For both strategies, we propose novel precoding and combining schemes. Specifically, we develop precoders based on the minimum mean square error (MMSE) criterion and investigate receive combining methods, including an MMSE combiner. Furthermore, we consider sharing different levels of channel knowledge between the APs. Finally, we present a comprehensive numerical analysis to validate our findings, evaluating the performance trade-offs associated with the number of users, APs, and antennas, as well as the choice of transmission strategy and the level of CSI sharing among APs.
@@ I Introduction
Cell-free massive multiple-input multiple-output (MIMO) is a wireless communication paradigm that has attracted great interest due to the vision of delivering uniformly high spectral efficiency (SE) over the coverage area [interdonato2019ubiquitous, chen2022survey, 9650567]. In cell-free massive MIMO, a large number of geographically distributed access points (APs) cooperate to serve the users on the same time-frequency resources through coherent joint transmission where interference is managed by MIMO methods [demir2021foundations].
@@
The recent surveys [9650567, chen2022survey] highlight a large body of work on cell-free massive MIMO systems. The predominant assumption is single-antenna users, although user devices (e.g., phones and tablets) have been equipped with multiple antennas since 4G. Four antennas are the current norm for mid-band 5G devices, and larger antenna numbers are expected in 6G devices operating at higher frequencies. The few papers on cell-free massive MIMO that consider multi-antenna users show SE improvements (e.g., [buzzi2017cell]) but assume perfect channel state information (CSI). A detailed SE analysis with realistic imperfect CSI is missing, particularly for downlink operation.
@@
The uplink SE with zero-forcing (ZF) combining and multi-antenna users was studied in [mai2019uplink] under the assumption of independent and identically distributed (i.i.d.) Rayleigh fading. This analysis was extended to consider the Weichselberger channel model in [wang2022uplink], where four operation regimes were compared. The uplink was further studied in [zhang2019performance, sun2023uplink] with a focus on distortion caused by low-resolution hardware. The downlink performance with ZF precoding is analyzed in [buzzi2019user] under i.i.d. Rayleigh fading, using the hardening bound where the user device lacks CSI. Moreover, the downlink performance with maximum ratio (MR) precoder and downlink channel estimation was studied in [mai2020downlink]. Furthermore, closed-form SE relations are given for the particular assumptions in their setting.While it is a foundational analysis of downlink SE in cell-free massive MIMO with multi-antenna users, its reliance on the assumption that the effective channel matrices converge to Gaussian matrices prevents generalization beyond simple precoders like the MR precoder. The downlink with low-resolution digital-to-analog converters (DACs) was considered in [zhou2021sum]. Multi-antenna users have also been studied in the context of cellular massive MIMO. A single-cell scenario was considered in [li2016massive] and the hardening bound was used to compute an achievable downlink SE without CSI at the users. [dovelos2020massive] computed a closed-form uplink SE expression in single-cell massive MIMO systems with multi-antenna users, MR combining, and Rician fading. Downlink MR precoding with multi-antenna users was considered in [sutton2021hardening], with a focus on making the hardening bound tighter with an unusual precoding normalization. [10571171] investigates the impact of spatial correlation and mutual coupling on SE. Whereas [wang2020uplink] analyzes uplink SE over spatially correlated Rician fading channels, comparing maximum ratio (MR) and local MMSE combining schemes. The correlation between the channel vectors to different antennas at the same user was studied in [wu2017novel], and it was shown that asymptotic orthogonality cannot be guaranteed. Synchronization issues, such as phase misalignments between APs, were studied in [ganesan2024cell] by proposing transmission methods and optimization of the AP clusters.
@@
In [ngo2017no], it was shown that no downlink pilots are needed in massive MIMO systems with single-antenna users. The receiving users can perform decoding properly without CSI since the effective channel’s phase is removed by precoding, and the gain approaches its mean value—a phenomenon called channel hardening. The so-called hardening bound on the capacity gives a performance close to the perfect CSI case also in cell-free massive MIMO [demir2021foundations] with the possible exception of scenarios with MR precoding, if the precoding vectors are normalized incorrectly [Chen2018a, Interdonato2021a] and when the effective channel means are normalized with their estimates [8799031]. In cases with very limited channel hardening (e.g., keyhole channels), higher SE can be achieved by explicitly estimating the effective channel’s amplitude from the received downlink signals [ngo2017no]. These single-antenna results might be the reason why the hardening bound has also been used in [buzzi2019user, zhou2021sum, li2016massive, sutton2021hardening] when considering multi-antenna users. However, we will show that the use of downlink pilots can lead to great SE improvements in these cases. [kanno2022fronthaul] proposes hybrid signal processing schemes for uplink cell-free massive MIMO, compressing received signals at APs to reduce fronthaul loads, while [shaik2025over] introduces an over-the-air (OTA) computation framework to minimize fronthaul overhead by computing global sufficient statistics locally at APs. Both these solutions are post-processing methods, whereas our work introduces a fundamentally different precoding method.
@@ I-A Motivation and Contributions
In this paper, we analyze the performance of a cell-free massive MIMO system with multiple antennas on the users and APs, in a scenario with correlated Rayleigh fading channels. Conventional cell-free massive MIMO builds on the assumption that the serving APs send the same data stream to the user. However, we can send multiple streams to multi-antenna users, which enables an alternative transmission method where the serving APs send different streams to the same user non-coherently. This separate-stream transmission mode allows each AP to forward an independent data stream to the UE, thereby eliminating the need for network-wide phase calibration and slashing the downlink fronthaul load. It represents an operating point that is fundamentally impossible with single-antenna UEs and has not been explored in earlier cell-free research. We analyze both transmission methods with arbitrary precoding to derive general SE expressions. In particular, we derive SE expressions for the cases when the user has CSI obtained through pilot-based downlink channel estimation. Unlike the scalar channels of single-antenna users, the effective MIMO channel seen by a multi-antenna UE is a matrix that does not harden to a deterministic scaled identity. Consequently, estimating this matrix via downlink pilots is indispensable, and our analysis shows that doing so closes the large spectral-efficiency gap between the hardening bound and the perfect-CSI benchmark.
We propose new precoders for the two transmission methods. These are needed because the conventional precoding schemes for multi-antenna users in [bandemer2006linear, stankovic2004multi, spencer2004zero] are not designed for situations with per-AP power constraints, partial CSI, and separate stream transmission. We further investigate different combining methods that suppress interference using the downlink channel estimates.
@@
In the conference version of this work [kama2024downlinkpilotsessentialcellfree], we considered a substantially simpler setup with i.i.d. Rayleigh fading, centralized operation, and ZF combining, which was sufficient to demonstrate the necessity of downlink pilots when having multi-antenna users. By contrast, in this paper, we modify the channel model to correlated Rayleigh fading (which requires more advanced estimation and more general SE expressions), explore three different transmission methods (two of which are novel contributions), and multiple receive combining methods (both MMSE and ZF combining). Moreover, we propose precoders tailored to the different transmission methods and derive novel SE relations for the general combiner case. These extensions and generalizations make the contributions applicable in practical conditions and provide insight into which design choices to make.
The rest of this paper is organized as follows. In Section II, we describe the channel model and estimation of the channel in the uplink. In Section II-B, downlink data transmission methods are explained for both same and separate stream methods. In Section III, achievable SE results for the same stream transmission method are derived, particularly with downlink pilots. In Section IV, SE results are extended to the case with separate stream transmission. Next, in Section V, precoders are proposed for all the considered transmission methods. We explain under which channel conditions these methods coincide. Section VII provides numerical results where we compare the performance of the considered methods and compare different SE expressions. Finally, conclusions and future directions are given in Section VIII.
[…]
@@ II System Model
We consider a cell-free massive MIMO system with $L$ APs and $K$ users arbitrarily distributed in a large geographical area. Each AP has $N$ antennas and each user has $M$ antennas. We consider the standard block-fading time-division duplexing (TDD) operation where the channel between AP $l$ and user $k$ in an arbitrary coherence block is denoted by $\mathbf{H}_{lk}\in\mathbb{C}^{N\times M}$ [demir2021foundations]. We assume a correlated Rayleigh fading channel model. Hence, the channel between AP $l$ and user $k$ is complex Gaussian distributed with the spatial correlation matrix $\mathbf{R}_{lk}\in\mathbb{C}^{NM\times NM}$, which is defined for the vectorization $\mathrm{vec}(\mathbf{H}_{lk})$ of the channel matrix $\mathbf{H}_{lk}$. The channel covariance is expressed through the covariances of its individual column vectors. To obtain this covariance, we vectorize the channel matrix.
@@ II-A Uplink Channel Estimation
As customary in TDD operation, the APs estimate the channels in the uplink and use the estimates in both uplink data reception and downlink data transmission [demir2021foundations]. In this paper, we will focus on uplink channel estimation and downlink data transmission. All users send uplink pilot sequences, and the APs estimate their channels to the users based on their received signals. Each user sends a $\tau_{\mathrm{p}}$-length orthonormal pilot sequence per transmit antenna. We will let $\tau_{\mathrm{c}}$ denote the length of the coherence block (in the number of symbols); thus, the pilot signaling overhead is $\tau_{\mathrm{p}}/\tau_{\mathrm{c}}$.
@@
$=\sqrt{q_{k}\tau_{\mathrm{p}}}\mathrm{vec}(\mathbf{H}_{lk})+\hskip-5.0pt\sum_{{i\in\mathcal{P}_{k}\setminus\{k\}}}\hskip-5.0pt\sqrt{q_{i}\tau_{\mathrm{p}}}\mathrm{vec}(\mathbf{H}_{li})+\mathrm{vec}(\mathbf{N}_{lk})$
@@
where we introduced the notation $\mathbf{h}_{li}=\mathrm{vec}(\mathbf{H}_{li})$ for the vectorization of the channel matrix and $\mathbf{n}_{lk}=\mathrm{vec}(\mathbf{N}_{lk})$ for the noise. The MMSE estimate of $\mathbf{h}_{lk}$ is [bjornson2017massive, Ch. 3]
@@ II-B1 Downlink with Same Stream Transmission
The most advanced downlink operation involves all APs in the transmission of all data streams, using the centralized operation [demir2021foundations]. Hence, the APs cooperate in the transmission, share CSI obtained from the uplink channel estimation, and precode their transmitted downlink signals based on it.
@@
For a given realization of the effective channels, created by the channel realizations and precoding scheme, the achievable SE varies depending on the available CSI at the receiver. We will provide expressions for three different cases in Section III and compare them numerically in Section VII.
@@ III-A Hardening Bound without CSI at the Receiver
When the receiver lacks CSI regarding the instantaneous channel realizations, the SE can be lower bounded using the hardening bound [demir2021foundations]. This method is conventionally used for single-antenna users, but the multi-antenna extension was considered in [buzzi2019user, zhou2021sum, li2016massive, mai2020downlink, wang2022uplink, sutton2021hardening]. The main idea is that user $k$ knows the statistics of its effective channel $\mathbf{B}_{kk}=\mathbf{H}_{k}^{\rm{H}}\mathbf{W}_{k}$, such as its mean $\bar{\mathbf{B}}_{kk}=\mathbb{E}\left\{\mathbf{B}_{kk}\right\}$, and the precoder matrix has been selected based on the CSI to make the variations around the mean small. We rewrite the received signal in (II-B1) as
@@
where $|\cdot|$ denotes the determinant.
The proof of this lemma builds on interpreting (8) as a point-to-point MIMO system with the deterministic channel matrix $\bar{\mathbf{B}}_{kk}$ and uncorrelated additive noise with the covariance matrix $\mathbf{\Xi}_{k}\in\mathbb{C}^{M\times M}$, which can be obtained by using sample averages. The SE expression is then obtained as a lower bound on the channel capacity by utilizing the worst-case uncorrelated additive noise term [Hassibi2003a]. The single-antenna version of Lemma 1 is called the hardening bound [demir2021foundations] since the effective channel realization $\mathbf{B}_{kk}$ is close to its mean $\bar{\mathbf{B}}_{kk}$ (in relative terms) when $M=1$, NN is large, and the precoder is selected as MR, ZF, or a filter of similar structure. This phenomenon is called channel hardening. However, as the effective channel is a matrix instead of a scalar when $M>1$, the same kind of channel hardening is not achieved herein since each column of the precoder matrix can only achieve hardening with respect to one row in the channel matrix. The resulting performance loss is demonstrated later in this paper.
@@
We conclude that only the diagonal entries of $\mathbf{B}_{kk}$ harden, while the off-diagonal entries that represent the inter-stream interference do not. There are no off-diagonal entries when $M=1$, so channel hardening holds in that special case (as shown in [ngo2017no, demir2021foundations]), while most entries do not harden when $M>1$. Therefore, the hardening bound is loose when $M>1$, which calls for developing a methodology for acquiring CSI at the receiving user using downlink pilots. [^1] The pilot-free downlink CSI acquisition method proposed in [ngo2017no] cannot be extended to the multi-antenna case because it relies on the proposal that the precoder matrix is selected to make the entries of $\mathbf{B}$ (approximately) real-valued and positive. This can only be guaranteed for $M$ entries of $\mathbf{B}$, while the remaining $M^{2}-M$ (typically off-diagonal entries) cannot be estimated in this way.